OpenStudy (anonymous):

An investor invested a total of $1,300 in two mutal funds. One fund earns a 5% profit while the other earned a 3% profit. If the investor's total profit was $55, how much was invested in each mutal fund?

OpenStudy (anonymous):

The total amount of money invested is $1300, so if "x" dollars go into the first fund, then "1300 - x" dollars go into the second fund.

OpenStudy (anonymous):

And you are given the total profit as $55, so let's write an equation that uses the profit information... the total profit and the amount of profit earned in each fund.

OpenStudy (anonymous):

Total profit = (amount in fund A)(profit percentage from fund A) + (amt in B)(profit % from B)

OpenStudy (anonymous):

Total profit = $55 = x * 5% + (1300 - x) * 3%

OpenStudy (anonymous):

see I am lost becsuse I keep multiplying it but getting 65 wihch I know is wrong, let me try again

OpenStudy (anonymous):

would it be 39.05?

OpenStudy (anonymous):

You need to solve for x...that's the amount in the 5% account. Then subtract that x amount from 1300 to get the amount in the 3% account.

OpenStudy (anonymous):

That is where I am confused at, I don't know how to solve for x

OpenStudy (anonymous):

55 = x * 5% + (1300 - x) * 3% To make it easier, why don't you multiply everything through by 100 to get rid of the percentages... 100 * 55 = 100 * x * 5% + 100 * (1300 - x) * 3% Simplifying (on the right side, 100 times the percentage leaves just 5 and 3): 5500 = 5x + (1300 - x)(3)

OpenStudy (anonymous):

So: 5500 = 5x + (1300 - x)(3) Then simplify the last part by distributing the 3: 5500 = 5x + (3900 - 3x) And simplify again: 5500 - 3900 = 2x Then: 1600 = 2x which means: 800 = x

OpenStudy (anonymous):

It's just a matter of step by step, moving terms around to isolate "x".

OpenStudy (anonymous):

I'm sorry, but I have to run... but I think you have good help arriving :)

OpenStudy (anonymous):

Thank you

hero (hero):

Total amount invested x + y = 1300 Total profit: .03x + .05y = 55 Solve using systems of equations: x + y = 1300 .03x + .05y = 55 .03x + .03y = 39 .03x + .05y = 55 .03x + .05y = 55 -.03x - .03y = -39 .02y = 16 y = 16/.02 y = 800 x = 500

OpenStudy (anonymous):

Okay so which one is 5% and 3%?

OpenStudy (anonymous):

lol ok